OpenStudy (anonymous):

Given v(s)=ds/dt=2s^2+3. Find a=d^2s/dt^2 as a function of s.

OpenStudy (anonymous):

4s?

OpenStudy (ajprincess):

hmm ya.

OpenStudy (anonymous):

so what if the question said: given v(t)=ds/dt=2t^2+3. Find a=d^2s/dt^2 as a function of t.

OpenStudy (anonymous):

i just want to clarify\[\frac{ds}{dt}=2s^2+3\]\[\frac{d^2s}{dt^2}=?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the thing that is throwing me off is the "as a function of s" or "as a function of t" part

OpenStudy (anonymous):

According to what you did, shouldn't the [2s ^{2}+3\] be squared?

OpenStudy (anonymous):

2s^2+3

OpenStudy (anonymous):

ahh wait its not clear

hartnn (hartnn):

i think u made a small error here: \[a=\frac{d^2s}{dt^2}=\frac{d}{dt}(\frac{ds}{dt})=\frac{d}{ds}(\frac{ds}{dt})\frac{dt}{ds}\]

OpenStudy (anonymous):

\[(2s ^{2}+3)^{2}\]

OpenStudy (anonymous):

\[a=[\frac{d}{ds}(\frac{ds}{dt})]\times\frac{ds}{dt}=[\frac{d}{ds}(2s^2+3)]\times (2s^2+3)=4s(2s^2+3)\]

OpenStudy (anonymous):

d/ds=4s, right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so 4s(2s^2+3)x(2s^2+3) equals what?

OpenStudy (anonymous):

shouldn't it equal 4s(2s^2+3)^2

OpenStudy (anonymous):

using chain rule\[a=\frac{d^2s}{dt^2}=\frac{d}{dt}(\frac{ds}{dt})=[\frac{d}{ds}(\frac{ds}{dt})]\times[\frac{ds}{dt}]\]

OpenStudy (anonymous):

\[a=[\frac{d}{ds}(\frac{ds}{dt})]\times[\frac{ds}{dt}]=[\frac{d}{ds}(2s^2+3)]\times(2s^2+3)\] \(=[2s^2+3]'(2s^2+3)=4s(2s^2+3)\)

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

sorry dude, totally was reading it wrong

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

thank you